OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (5, 3) with a slope of -3/4.

OpenStudy (anonymous):

y=mx+c, where slope,m=-3/4,y=(-3/4)x+c,4y+3x=4c (5,3), 12+15=4c=27,c=27/4 (x,6) ,24+3x=4c 4c-24=3x=27-24=3 (by substituting c) so3x=3, x=1